OpenStudy (anonymous):

An automated car wash can wash 30 cars in 5 hours. How long would it take to wash 120 cars? A. 15 h B. 20 h C. 25 h D. 30 h

OpenStudy (anonymous):

help @siblings

OpenStudy (anonymous):

@SageWilson

OpenStudy (sagewilson):

it would take 20 hours, so B.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

i fanned you

OpenStudy (sagewilson):

thank you

OpenStudy (er.mohd.amir):

wash one car in =5/30 then 120 cars =120*(5/30)=600/20=30 h